OpenStudy (anonymous):

Which of the following points lies on the circle whose center is at the origin and whose radius is 5? A.) (-3, 4) B.) (1, -2) C.) (√5,√5 )

OpenStudy (anonymous):

the equation of a circle centered at the origin with radius of r is \[x^2+y^2=r^2\] This circle's equation is \[x^2+y^2=5\] Which of the points makes that equation true?

OpenStudy (anonymous):

you actually have to plug the numbers in to find the answer

OpenStudy (anonymous):

i think i am doing it wrong

OpenStudy (anonymous):

what did you get for x² + y² when you put in (-3, 4)?

OpenStudy (anonymous):



OpenStudy (anonymous):

25

OpenStudy (anonymous):

ok. that's not 5, so (-3,4) isn't the answer. Try (1, -2)

OpenStudy (anonymous):

5

OpenStudy (anonymous):

right. so (1, -2) is the answer since it gives 5 on the left to match the 5 on the right